OpenStudy (anonymous):

Please help: will medal or fan... In Mrs. Williams class there are 13 students. They each have 4 pieces of white paper. Some of the students, s, also have 3 pieces of blue paper each. In all, the students have 64 pieces of paper. Which equation represents this situation? (s × 3) + 13 = 64 s × 3 = 64 (13 × 3) + s = 64 (13 × 4) + (s × 3) = 64

OpenStudy (anonymous):

@bobsue @whpalmer4 @SectorZ423

OpenStudy (whpalmer4):

How many students are there?

OpenStudy (anonymous):

13?

OpenStudy (whpalmer4):

You tell me.

OpenStudy (anonymous):

yeah theirs 13

OpenStudy (whpalmer4):

"Theirs" is not a word in the English language. "There's" is a contraction of "there is", and would be the right choice here, except that as the number of students is more than one, it should be "there are". 13 students. How many pieces of white paper per student?

OpenStudy (whpalmer4):

Sorry, of course "theirs" is a word, just not a word that can be used as you did.

OpenStudy (whpalmer4):

I should be pickier about how I'm picky :-)

OpenStudy (anonymous):

only 4 are of each white and 3 of the blue were given to each student

OpenStudy (whpalmer4):

I'll try my question again, if you would be so kind as to try to answer it again. 13 students. How many pieces of white paper per student?

OpenStudy (anonymous):

4

OpenStudy (whpalmer4):

Thank you. If there are 13 students, and 4 pieces of white paper per student, what would be an expression for how many pieces of white paper there are?

OpenStudy (anonymous):

13 × 3) + s = 64

OpenStudy (anonymous):

13/4

OpenStudy (anonymous):

@myusernamesucks that was the wrong answer I'm doing studyisland and It has a red mrk next to it

OpenStudy (anonymous):

oops mark

OpenStudy (whpalmer4):

Why would you divide 13 by 4? 1 student has 4 pieces of white paper. 2 students have 4 pieces of white paper each, or 4+4 or 2*4 3 students have 4 pieces of white paper each, or 4+4+4 or 3*4 ... 13 students have 4 pieces of white paper each. The expression is?

OpenStudy (anonymous):

13*4=52 but I don't get it why would you multipy?

OpenStudy (whpalmer4):

Another way of thinking about this: \[\frac{4\text{ pieces of white paper}}{1 \text{ student}} * 13\text{ students} = \frac{4\text{ pieces of white paper}}{1 \cancel{\text{ student}}} * 13\cancel{\text{ students}} \]\[\qquad= \text{_____ pieces of white paper} \]

OpenStudy (whpalmer4):

If you have to give 3 pieces of gum to each of your 4 friends, how many pieces of gum do you need?

OpenStudy (anonymous):

oh... you could have said that the first time

OpenStudy (anonymous):

12 I like gum!!

OpenStudy (anonymous):

I did 4*3 and got 12

OpenStudy (whpalmer4):

And how did you come up with that answer? By multiplying, right? Whether you realize it or not, you did that calculation with the fractions that I wrote out. Okay, I asked for the expression, not the number. What will be the expression for 13 students each having 4 pieces of white paper?

OpenStudy (whpalmer4):

We are trying to find the pieces that will be assembled to be our answer. 52 doesn't appear in any of the answers.

OpenStudy (anonymous):

yeah of course and I found my answer while you were typing... thanks snyway

OpenStudy (whpalmer4):

I suggest you continue going through it with me, so that you develop some proficiency at analyzing these situations and writing the correct expression.

OpenStudy (whpalmer4):

Ah, it seems you're more interested in moving on to the next problem. Perhaps if you'd stuck around and finished listening to what I had to say about the problem here, you wouldn't need help with the next problem. Good luck!